# Soft spots of net negative topological charge directly cause the plasticity of 3D glasses

### v2

@@ Abstract
We examine the spatial distribution and probability distribution of connected cells that had high net charge imbalance of the same sign. We then calculate the normalized cluster size distribution $P(n_{c})$, where $n_{c}$ represents the number of connected cubical cells with a high topological charge of the same sign. In Fig.6(a)-(d), we depict the cluster size distribution $P(n_{c})$ for $\bar{C}^{c}_{k}>0.6$ and $\bar{C}^{c}_{k}<-0.6$ during yielding. We have highlighted the three largest clusters in the distributions for $\gamma=0.098$ and $\gamma=0.099$ in Fig. 6(a)-(b). Right before yielding, we observe the emergence of very large clusters, with the largest one being of negative net charge imbalance. Right after yielding, these large clusters disappear and the distribution tends to narrow and fragment into clusters of smaller sizes. The rapid formation and disappearance of large clusters of negative topological charge – a “topological catastrophe” – suggests that yielding might be caused by the massive annhilation of topological defects. Furthermore, following the recent results presented in Desmarchelier et al. (2024), the large clusters of negative topological charge, related to STZ, appearing before yielding might be a manifestation of the formation of shear bands. It is also plausible that the structure and dynamics of these defects bear strong connections with the concept of elastic screening discussed in Lemaître et al. (2021).
@@
Conclusion – In this study, we systematically explored the formation, dynamics and statistical properties of topological defects in a 3D polymer glass model under shear deformation. We first established that, in 3D glasses, the topological defects with -1 charge (saddles or anti-vortices) are associated with large non-affine displacements and act as triggers of local plastic activity. Also, the total defects density anti-correlates with the non-affine displacements magnitude as a function of shear strain. This suggests that the transient accumulation of defects triggers a subsequent plastic event, upon further increasing the strain, via massive annihilation of negative and positive defects. Overall, the deformation process upon increasing the strain is characterized by a dominance of negative charges over positive charges, which is linked to the tendency of the polymer glass to strain-soften prior and across yielding. Through a coarse-graining procedure into mesoscopic cells, we unveiled the formation of very few, but very large, clusters bearing net negative topological charge right before the yielding point. These clusters are associated with huge values of the non-affine displacements magnitude. The emergence of these very large negative clusters right before yielding is followed by its massive annihilation right at yielding, which is accompanied by a system-spanning plastic event marking the onset of plastic flow.
@@
Several open questions remain to be answered. As already anticipated in Desmarchelier et al. (2024), topological defects with negative charge exhibit strong resemblances with quasi-localized quadrupolar fields previously identified as the carriers of plasticity in amorphous solids Dasgupta et al. (2012); Hieronymus-Schmidt et al. (2017). Indeed, topological defects can be identified as geometric singularities even in amorphous systems Moshe et al. (2015); Livne et al. (2023), and might constitute the building blocks for the dipolar and quadrupolar objects discussed in the literature in relation to elastic screening and anomalous elasticity Kumar and Procaccia (2023). These topological defects could provide the missing link to understand yielding in amorphous systems as a topological phase transition Livne et al. (2023); Jin et al. (2024), akin to the famous BKT transition in 2D crystals Zippelius et al. (1980); Fraggedakis et al. (2023).
@@
As a final remark, despite what has been thought for long time, topological concepts are still valid beyond the realm of structural order, and they emerge as the best candidates to characterize plasticity and mechanical failure in amorphous materials from a microscopic point of view, as achieved for the crystalline counterparts.
Methods – We use the Kremer-Grest model Kremer and Grest (1986) to investigate a coarse-grained polymer system consisting of linear chains, each composed of $50$ monomers. The monomer masses within a polymer chain alternate between two values, $m_{1}=1$ and $m_{2}=3$, with a total of $N=10,000$ monomers in the system.
@@
Acknowledgments – We would like to thank Y. Jin, D. Pan, W. Zhang, J. Zhang, M. Pica Ciamarra, M. Moshe and C. Jiang for useful discussions and collaboration on related topics. M.B. acknowledge the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01). MB acknowledges the support of the sponsorship from the Yangyang Development Fund. ACYL acknowledges support from the Australian Research Council (FT180100594). A.Z. gratefully acknowledges funding from the European Union through Horizon Europe ERC Grant number: 101043968 “Multimech”, and from the Niedersächsische Akademie der Wissenschaften zu Göttingen in the frame of the Gauss Professorship program. A.Z. and A.B. gratefully acknowledge funding from US Army Research Office through contract nr. W911NF-22-2-0256.

### v3

@@ Results
We examine the spatial distribution and probability distribution of connected cells that had high net charge imbalance of the same sign. We then calculate the normalized cluster size distribution $P(n_{c})$, where $n_{c}$ represents the number of connected cubical cells with a high topological charge of the same sign. In Fig.8(a)-(d), we depict the cluster size distribution $P(n_{c})$ for $\bar{C}^{c}_{k}>0.6$ and $\bar{C}^{c}_{k}<-0.6$ during yielding. We have highlighted the three largest clusters in the distributions for $\gamma=0.098$ and $\gamma=0.099$ in Fig. 8(a)-(b). Right before yielding, we observe the emergence of several large clusters, with the largest one being of negative net charge imbalance. At yielding, many of these large clusters disappear and the distribution tends to narrow and fragment into clusters of smaller sizes. After each PE, the system again tends to develop negative charged clusters (see table 1). The rapid formation and disappearance of large clusters of negative topological charge – a “topological avalanche” Altshuler and Johansen (2004)– suggests that yielding might be caused by the massive annhilation of topological defects. Furthermore, following the recent results presented in Desmarchelier et al. (2024), the large clusters of negative topological charge, related to STZ, appearing before yielding might be a manifestation of the formation of shear bands. It is also plausible that the structure and dynamics of these defects bear strong connections with the concept of elastic screening discussed in Lemaître et al. (2021).
Figure 7: The simulation boxes contain cubical cells with negative charge ($\bar{C}^{c}_{k}<-0.6$) shaded in light blue. Within this representation, we present clusters with size $n_{c}\geq 8$ from this negative charge cloud. (a) At $\gamma=0.098$ (just before yielding), 11 clusters are displayed, colored in red ($n_{c}=17$, $N_{n_{c}}=1$), green ($n_{c}=15$, $N_{n_{c}}=1$), magenta ($n_{c}=14$, $N_{n_{c}}=1$), yellow ($n_{c}=10$, $N_{n_{c}}=1$), cyan ($n_{c}=9$, $N_{n_{c}}=3$), and purple ($n_{c}=8$, $N_{n_{c}}=4$), respectively. (b) At $\gamma=0.099$ (at yielding), the 6 largest clusters are shown in red ($n_{c}=31$, $N_{n_{c}}=1$), green ($n_{c}=13$, $N_{n_{c}}=1$), magenta ($n_{c}=10$, $N_{n_{c}}=1$), and yellow ($n_{c}=9$, $N_{n_{c}}=3$). (c) At $\gamma=0.153$ (just before another PE), 11 largest clusters are presented in red ($n_{c}=21$, $N_{n_{c}}=1$), green ($n_{c}=17$, $N_{n_{c}}=1$), magenta ($n_{c}=16$, $N_{n_{c}}=1$), yellow ($n_{c}=12$, $N_{n_{c}}=4$), cyan ($n_{c}=11$, $N_{n_{c}}=2$), and purple ($n_{c}=8$, $N_{n_{c}}=2$). (d) At $\gamma=0.154$ (occurrence of another PE), 3 clusters are displayed in red ($n_{c}=10$, $N_{n_{c}}=1$), green ($n_{c}=9$, $N_{n_{c}}=1$), magenta ($n_{c}=8$, $N_{n_{c}}=1$), respectively. Note that $N_{n_{c}}$ is the number of clusters of size $n_{c}$.
@@ Conclusion
In this study, we systematically explored the formation, dynamics and statistical properties of topological defects in a 3D polymer glass model under shear deformation. We first established that, in 3D glasses, the topological defects with -1 charge (saddles or anti-vortices) are associated with large non-affine displacements and act as triggers of local plastic activity. Also, the total defects density anti-correlates with the non-affine displacements magnitude as a function of shear strain. This suggests that the transient accumulation of defects triggers a subsequent plastic event, upon further increasing the strain, via massive annihilation of negative and positive defects. Overall, the deformation process upon increasing the strain is characterized by a dominance of negative charges over positive charges, which is linked to the tendency of the polymer glass to strain-soften prior and across yielding. Through a coarse-graining procedure into mesoscopic cells, we unveiled the formation of several large clusters bearing net negative topological charge before the yielding point. These clusters are associated with huge values of the non-affine displacement magnitude. The emergence of these large negative clusters before yielding is followed by its massive annihilation right at yielding, which is accompanied by a system-spanning plastic event marking the onset of plastic flow. While our findings pertain to polymer glasses, we expect the same methodology to yield interesting insights into the deformation behaviour of other amorphous and glassy systems.
@@
Several open questions remain to be answered. As already anticipated in Desmarchelier et al. (2024), topological defects with negative charge exhibit strong quasi-localized quadrupolar field previously identified as the carriers of plasticity in amorphous solids Dasgupta et al. (2012); Hieronymus-Schmidt et al. (2017). Indeed, topological defects can be identified as geometric singularities even in amorphous systems Moshe et al. (2015); Livne et al. (2023), and might constitute the building blocks for the dipolar and quadrupolar field discussed in the literature in relation to elastic screening and anomalous elasticity Kumar and Procaccia (2023). These topological defects could provide the missing link to understand yielding in amorphous systems as a topological phase transition Livne et al. (2023); Jin et al. (2024), akin to the famous Berezinskii-Kosterlitz-Thouless (BKT) transition in 2D crystals Zippelius et al. (1980); Fraggedakis et al. (2023). In our study, we characterized the topological defects by analyzing quasi-2D slices of the three-dimensional system. It would be interesting to investigate 3D defects within the fully three-dimensional non-affine field. This direct 3D characterization of topological defects in glasses has been less explored in the literature (see nevertheless Cao et al. (2018)) and is computationally more challenging. This presents an intriguing avenue for future research.
Vortex-like topological defects have been originally proposed in the eigenvector field Wu et al. (2023), while our study generalizes this concept to the displacement vector field. At this point, it is not entirely clear which are the differences and similarities between the two approaches. Nevertheless, following the aforementioned results of Tanguy et al. (2002) it is plausible to expect a correlation between the topological features in the displacement field and those in the low-frequency eigenvector fields around strong plastic events. The main reason behind this expected correlation is the fact that plastic events are typically dominated by low-frequency modes that correspond to the disappearance of local potential barriers in the potential energy landscape. A more detailed comparison between these two methods is left for future research. See also Desmarchelier et al. (2024) for an analysis of topological defects in the displacement vector field.
@@
As a final remark, despite what has been thought for long time, topological concepts are still valid beyond the realm of structural order. Here we show topological defects emerge as extremely promising candidate to characterize plasticity and mechanical failure in amorphous materials at a microscopic level, as achieved for their crystalline counterparts.
Appendix: Simulation methods
We use the Kremer-Grest model Kremer and Grest (1986) to investigate a coarse-grained polymer system consisting of linear chains, each composed of $50$ monomers. The monomer masses within a polymer chain alternate between two values, $m_{1}=1$ and $m_{2}=3$, with a total of $N=10,000$ monomers in the system.
@@
Supplementary Material
Supplementary material is available at PNAS Nexus online.
Data availability
All data for this study are presented in the manuscript and the Supplementary material. Additional data (polymer glass models and interpolated fields) are available from the Figshare repository: DOI: 10.26180/26199917.
Funding
M.B. acknowledge the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01). MB acknowledges the support of the sponsorship from the Yangyang Development Fund. ACYL acknowledges support from the Australian Research Council (FT180100594). A.Z. gratefully acknowledges funding from the European Union through Horizon Europe ERC Grant number: 101043968 “Multimech”, and from the Niedersächsische Akademie der Wissenschaften zu Göttingen in the frame of the Gauss Professorship program. A.Z. and A.B. gratefully acknowledge funding from US Army Research Office through contract nr. W911NF-22-2-0256.
Author contributions statement
A. Bera performed the numerical simulations and data analysis with the help of A. Liu. A. Bera wrote the manuscript with the help of the other authors. A. Zaccone supervised the project. All authors contributed to the development of the ideas behind this work and the physical interpretation of the results.
Acknowledgements.
We would like to thank Y. Jin, D. Pan, W. Zhang, J. Zhang, M. Pica Ciamarra, M. Moshe and C. Jiang for useful discussions and collaboration on related topics.